# FLLIC: Functionally Lossless Image Compression

### v3

@@ IV-A Experiment Setup
These cascaded methods are evaluated to understand the impact of different denoising and compression combinations on the overall performance.
@@
These lossy compression methods serve as strong baselines to demonstrate how FLLIC performs in comparison to advanced lossy codecs that do not specifically target noise removal.
@@
For a fair comparison, we fine-tune all baselines on the Flickr2K dataset. Pre-trained weights for each chosen method are used as initialization, and the entire fine-tuning process is conducted under conditions similar to those used for our FLLIC method. This ensures that all baselines are well-adapted to the noisy or near-lossless input scenarios, allowing for a rigorous and meaningful evaluation.
@@ IV-B Results on Synthetic Datasets
Based on the results shown in Fig. 3, we can draw several key observations regarding the rate-distortion (R-D) performance of various compression methods across different noise levels ($\sigma=1,2,3$) on the Kodak and DIV2K-val datasets.
The proposed FLLIC method consistently achieves the highest PSNR (Peak Signal-to-Noise Ratio) at any given BPSP (bits per sub-pixel) across all noise levels ($\sigma=1,2,3$). It consistently outperforms all other methods in both the Kodak and DIV2K-val datasets, demonstrating superior joint denoising and compression capabilities. FLLIC is especially effective in higher noise scenarios ($\sigma=2$ and $3$), where it maintains better compression efficiency and higher PSNR compared to other methods, reflecting its ability to effectively remove noise while maintaining high fidelity in the reconstruction.
@@
In conclusion, the FLLIC method provides the best performance across a range of noise levels ($\sigma=1,2,3$) and datasets (Kodak and DIV2K-val). Its ability to simultaneously denoise and compress images leads to superior PSNR values compared to all other baselines, including lossy and near-lossless methods. The weakly-supervised FLLIC-w offers a strong alternative when fully supervised data is unavailable. The comparison with other denoising and compression baselines shows that FLLIC’s joint denoising and compression strategy delivers a more efficient and accurate solution, especially in higher-noise scenarios.
@@ IV-C Results on Real-world Dataset
Fig. 4 presents the Rate-Distortion (R-D) performance comparison of various methods on the real-world SIDD dataset. The results demonstrate the efficacy of the proposed FLLIC method in comparison to a variety of denoising and compression approaches under realistic noise conditions. The proposed FLLIC method stands out by achieving the highest PSNR values across all BPSP levels. Notably, it outperforms all other methods, including lossy compression methods (ELIC and MLIC), as well as cascaded denoising + lossless compression techniques. The results highlight FLLIC’s exceptional ability to simultaneously denoise and compress images, effectively mitigating real-world noise in the process.
FLLIC-w, the weakly-supervised version of FLLIC, follows closely behind but shows a slight decrease in performance compared to the fully-supervised FLLIC. This indicates that the weak supervision does not significantly degrade the performance, especially at higher BPSP values, making FLLIC-w a viable option when fully supervised training data is unavailable.
The Restormer + LC-FDNet approach, which combines transformer-based denoising with lossless compression, demonstrates competitive performance in comparison to FLLIC, particularly at lower BPSP values. However, it still lags behind FLLIC in higher BPSP regions. Similarly, CGNet + ArIB-BPS performs well but struggles to match the performance of FLLIC in this dataset. These findings suggest that while cascaded denoising and compression methods can offer promising results, they are limited by their ability to jointly optimize denoising and compression simultaneously.
In terms of lossy compression, both MLIC and ELIC perform notably worse than FLLIC, especially at higher BPSP values. MLIC and ELIC are optimized for lossy compression, but they do not leverage the denoising aspect that is crucial for real-world noisy images. MLIC, while improving PSNR over traditional lossy compression methods, still falls behind FLLIC due to its inability to handle noise as effectively. Similarly, ELIC, despite being a learned lossy compression method, does not achieve the level of noise suppression and image fidelity that FLLIC provides, making it less suitable for real-world denoising and compression tasks.
In summary, the FLLIC method exhibits superior rate-distortion performance on the real-world SIDD dataset, confirming its robustness and effectiveness in joint denoising and compression. The results further emphasize the advantages of FLLIC over both traditional and learned methods, particularly in the context of real-world noisy image compression tasks.es.
Figure 5: Rate-Distortion performance comparison of FLLIC, FLLIC-w, and DLPR (near-lossless compression) methods on Kodak dataset with varying noise levels ($\sigma=1,2,3$). The x-axis represents the bits per sub-pixel (BPSP), and the y-axis denotes the peak signal-to-noise ratio (PSNR) in dB. The plots show the performance of FLLIC and its weakly-supervised version (FLLIC-w) alongside the near-lossless compression method DLPR for three different values of $\tau$.
@@ IV-D Comparison with Near-lossless Compression
Fig. 5 shows the Rate-Distortion (R-D) performance comparison of FLLIC, FLLIC-w, and DLPR (near-lossless compression) methods on the Kodak dataset, with varying noise levels ($\sigma=1,2,3$). The plots compare the performance of FLLIC and its weakly-supervised version (FLLIC-w) with the near-lossless compression method DLPR for three different values of $\tau$.
At noise level $\sigma=1$ (first plot), FLLIC achieves the highest PSNR ( 49 dB) at a BPSP of around 3.0, with FLLIC-w performing slightly lower ( 48.5 dB). The near-lossless compression method DLPR at $\tau=1$ shows a performance drop ( 47.8 dB), and as $\tau$ increases, its performance further declines, with $\tau=3$ reaching the lowest PSNR ( 45 dB). At $\sigma=2$ (second plot), the performance of all methods decreases as noise increases, but FLLIC still outperforms the others, achieving PSNR of around 46.5 dB at a BPSP of 2.8. DLPR’s performance at $\tau=1$ continues to drop ( 44 dB), and the gap between FLLIC and DLPR widens as $\tau$ increases. At $\sigma=3$ (third plot), FLLIC remains the best-performing method ( 45 dB), while DLPR’s PSNR continues to degrade, reaching as low as 40 dB at $\tau=3$. These results demonstrate FLLIC’s superior performance and robustness, especially in low-noise scenarios and at higher noise levels, compared to DLPR.
In all noise scenarios, FLLIC provides the best performance in terms of PSNR, followed by FLLIC-w. The performance of DLPR significantly degrades as $\tau$ increases, particularly in the presence of noise. This indicates that FLLIC and its weakly-supervised version are more effective at denoising and compression across different noise levels, while DLPR is more sensitive to noise and the value of $\tau$. FLLIC’s consistent performance underscores its robustness and effectiveness for real-world image compression tasks, especially under noisy conditions.
@@
Cascaded Cascaded Lossy Near-lossless Proposed FLLIC
@@ IV-E Inference Time
IV-F Limitation
The main limitation of this work is its need to train a specific network for a given noise level. The current version hardly generalizes to various noise levels by a single network. A possible solution could be estimating noise level first and using the estimated noise level as known a prior for the encoding and decoding. We leave detailed study of this limitation to the further work.

### v4

@@ IV-A Experiment Setup
For a fair comparison, we fine-tune all baselines on the Flickr2K dataset. Pre-trained weights for each chosen method are used as initialization, and the entire fine-tuning process is conducted under conditions similar to those used for our FLLIC method.
@@ IV-B Results on Synthetic Datasets
Based on the results shown in Figure 4, we can draw several key observations regarding the rate-distortion (R-D) performance of various compression methods across different noise levels ($\sigma=1,2,3$) on the Kodak and DIV2K-val datasets. The proposed FLLIC method consistently achieves the highest PSNR (Peak Signal-to-Noise Ratio) at any given BPSP (bits per sub-pixel) across all noise levels ($\sigma=1,2,3$). It consistently outperforms all other methods in both the Kodak and DIV2K-val datasets, demonstrating superior joint denoising and compression capabilities. FLLIC is especially effective in higher noise scenarios ($\sigma=2$ and $3$), where it maintains better compression efficiency and higher PSNR compared to other methods, reflecting its ability to effectively remove noise while maintaining high fidelity in the reconstruction.
@@
The MLIC and ELIC methods, both based on learned lossy compression techniques, originally performed significantly worse than the lossless and near-lossless methods (such as FLLIC and FLLIC-w) across all noise levels. This is expected, as standard lossy compression methods inherently prioritize rate–distortion trade-offs with respect to the noisy input, often sacrificing reconstruction fidelity to improve bitrate. To ensure a fairer comparison under the functionally lossless evaluation protocol, we retrained both MLIC and ELIC using noisy–clean image pairs, aligning their objectives with those of FLLIC. The retrained variants, denoted as MLIC and ELIC, show noticeable improvements in PSNR compared to their original versions. However, despite these improvements, both retrained variants still fall far behind FLLIC in terms of rate–distortion performance, requiring substantially higher bitrates yet failing to match FLLIC’s reconstruction quality.
@@ IV-C Results on Real-world Dataset
Figure 5 presents the Rate-Distortion (R-D) performance comparison of various methods on the real-world SIDD dataset. The results demonstrate the efficacy of the proposed FLLIC method in comparison to a variety of denoising and compression approaches under realistic noise conditions. The proposed FLLIC method stands out by achieving the highest PSNR values across all BPSP levels. Notably, it outperforms all other methods, including lossy compression methods (ELIC and MLIC), as well as cascaded denoising + lossless compression techniques. The results highlight FLLIC’s exceptional ability to simultaneously denoise and compress images, effectively mitigating real-world noise in the process.
FLLIC-w, the weakly-supervised version of FLLIC, follows closely behind but shows a slight decrease in performance compared to the fully-supervised FLLIC. This indicates that the weak supervision does not significantly degrade the performance, especially at higher BPSP values, making FLLIC-w a viable option when fully supervised training data is unavailable. The Restormer + LC-FDNet approach, which combines transformer-based denoising with lossless compression, demonstrates competitive performance in comparison to FLLIC, particularly at lower BPSP values. However, it still lags behind FLLIC in higher BPSP regions. Similarly, CGNet + ArIB-BPS performs well but struggles to match the performance of FLLIC in this dataset. These findings suggest that while cascaded denoising and compression methods can offer promising results, they are limited by their ability to jointly optimize denoising and compression simultaneously.
In terms of lossy compression, both MLIC and ELIC perform notably worse than FLLIC, especially at higher BPSP values. MLIC and ELIC are optimized for lossy compression, but they do not leverage the denoising aspect that is crucial for real-world noisy images. MLIC and ELIC, while improving PSNR over traditional lossy compression methods, still fall behind FLLIC due to its inability to handle noise as effectively.
Figure 6: Rate-Distortion performance comparison of FLLIC, FLLIC-w, and DLPR (near-lossless) methods on Kodak dataset with varying noise levels ($\sigma=1,2,3$). The x-axis represents the bits per sub-pixel (BPSP), and the y-axis denotes the peak signal-to-noise ratio (PSNR) in dB. The plots show the performance of FLLIC and its weakly-supervised version (FLLIC-w) alongside the near-lossless compression method DLPR for three different values of $\tau$.
@@ IV-D Comparison with Near-lossless Compression
Figure 6 shows the Rate-Distortion (R-D) performance comparison of FLLIC, FLLIC-w, and DLPR (near-lossless compression) methods on the Kodak dataset, with varying noise levels ($\sigma=1,2,3$). The plots compare the performance of FLLIC and its weakly-supervised version (FLLIC-w) with the near-lossless compression method DLPR for three different values of noise level $\tau$.
In all noise scenarios, FLLIC provides the best performance in terms of PSNR, followed by FLLIC-w. Specifically, at noise level $\sigma=1$ (first plot), FLLIC achieves the highest PSNR ( 49 dB) at a BPSP of around 3.0, with FLLIC-w performing slightly lower ( 48.5 dB). The near-lossless compression method DLPR at $\tau=1$ shows a performance drop ( 47.8 dB), and as $\tau$ increases, its performance further declines, with $\tau=3$ reaching the lowest PSNR ( 45 dB). At $\sigma=2$ (second plot), the performance of all methods decreases as noise increases, but FLLIC still outperforms the others, achieving PSNR of around 46.5 dB at a BPSP of 2.8. DLPR’s performance at $\tau=1$ continues to drop ( 44 dB), and the gap between FLLIC and DLPR widens as $\tau$ increases. At $\sigma=3$ (third plot), FLLIC remains the best-performing method ( 45 dB), while DLPR’s PSNR continues to degrade, reaching as low as 40 dB at $\tau=3$. These results demonstrate FLLIC’s superior performance and robustness, especially in low-noise scenarios and at higher noise levels, compared to DLPR.
@@
Cascaded Lossy Near-lossless Proposed FLLIC
@@ IV-E Inference Time
IV-F Visualization of Noise-Adaptive Quantization Maps
Understanding how noise statistics influence the learned content-adaptive quantization maps is important for interpreting the behavior of our proposed FLLIC framework. In this subsection, we provide both a rationale for the design and empirical visualizations to clarify this mechanism.
Design rationale.
It is intractable to derive an explicit, closed-form mapping from local noise characteristics to optimal quantization step sizes. First, the relevant statistics are determined jointly by the underlying clean signal and the spatially varying noise power. Second, the distributions of latent and hyper-latent representations after nonlinear analysis transforms are highly signal-dependent. To overcome these challenges, FLLIC leverages the estimated clean entropy $H(I)$ as an informative proxy that implicitly captures the effect of noise on local compressibility. By conditioning the quantization step prediction networks on $H(I)$, the model learns to allocate larger quantization steps in regions with higher noise and finer steps in regions rich in structural content. This adaptive behavior emerges automatically through end-to-end training and does not require handcrafted analytical formulas.
Visualization results.
To provide empirical evidence of this learned behavior, we visualize the predicted $Q_{y}$ (for main latents) and $Q_{z}$ (for hyper-latents) maps under different global noise levels. Specifically, we use the same test image corrupted with uniform Gaussian noise at three standard deviations $\sigma=1,2,3$, and plot the corresponding heatmaps of the predicted quantization steps. Figure 7 shows two key trends. First, as the global noise level increases, the quantization steps overall become larger, reflecting the reduced need to preserve fine details at lower signal-to-noise ratios. Second, within each noise level, the maps consistently assign finer steps along structural edges and coarser steps in smoother areas, indicating that the model also adapts to local image content. These observations confirm that our learned quantization maps respond simultaneously to global noise statistics and local structural importance.
Figure 7: Visualization of learned quantization maps ($Q_{y}$ and $Q_{z}$) for four example images under varying noise levels. Each row corresponds to a different input image, and within each, quantization step sizes increase with global noise level while retaining finer steps near structural edges.
IV-G Factorized vs. Auto-Regressive Entropy Models
In our framework, the entropy model for the hyperlatents $z$ adopts a fully factorized form. This design choice is widely adopted in learned image compression due to the typically minor contribution of hyperlatents to the overall bitrate and the desire to maintain fast, parallel decoding. In contrast, complex context models such as auto-regressive masked convolutions can offer marginal compression gains at the cost of significantly increased computational overhead.
To empirically validate this design choice, we additionally implemented an auto-regressive variant of FLLIC (denoted as FLLIC-A) that employs masked convolutions to capture spatial dependencies within $z$. Both versions were evaluated on the Kodak dataset under AWGN noise levels $\sigma=1,2,3$. The results are summarized in Table II.
TABLE II: Comparison between factorized (FLLIC) and auto-regressive (FLLIC-A) entropy models on Kodak. Reported: BPP / PSNR (dB) / Inference time change.
$\sigma$ FLLIC FLLIC-A $\Delta$ BPP $\Delta$ PSNR $\Delta$ Time
1 3.009 / 48.91 2.960 / 49.03 $-1.6\%$ $+0.12$ $+105\%$
2 2.959 / 45.71 2.915 / 45.82 $-1.5\%$ $+0.11$ $+113\%$
3 2.802 / 43.88 2.760 / 44.04 $-1.5\%$ $+0.16$ $+120\%$
As shown, the auto-regressive model provides only marginal bitrate reduction ($\sim\!1.5\%$) and minor PSNR gains ($<0.2$ dB), while incurring over $100\%$ increase in inference latency. Considering the negligible compression improvement but substantial runtime penalty, we retain the factorized entropy model as the default for its superior balance between efficiency and performance in practical deployment.
IV-H Ablation on Entropy Estimation and Adaptive Quantization
To evaluate the contribution of key modules in FLLIC, we conduct an ablation study on the Kodak dataset with synthetic AWGN noise ($\sigma=1,2,3$). Specifically, we progressively disable the entropy estimation branch $H(I)$ and the content-adaptive quantization modules $Q_{y},Q_{z}$. The results are tabulated in Table III and discussed below: (i) Removing $H(I)$ and replacing it with a global scalar increases the bitrate by $6.8\%$ on average and causes a slight but consistent PSNR drop ($\leq 0.1$ dB). This confirms that entropy estimation provides a strong global prior to reduce redundancy. (ii) Disabling spatially adaptive $Q_{y},Q_{z}$ further increases bitrate by $9.2\%$ and leads to slightly larger PSNR degradation ($\sim\!0.1$ dB), indicating that local adaptation is essential for handling heterogeneous content. (iii) Removing both components causes over $15\%$ bitrate deterioration and up to $0.18$ dB loss in PSNR, highlighting their complementary roles.
TABLE III: Ablation study on entropy estimation and adaptive quantization. Reported: BPP / PSNR (dB).
$\sigma$ FLLIC (full) w/o $H(I)$ w/o $Q_{y},Q_{z}$ w/o Both
1 3.009 / 48.91 3.214 / 48.85 3.285 / 48.82 3.476 / 48.77
2 2.959 / 45.71 3.160 / 45.66 3.230 / 45.62 3.417 / 45.55
3 2.802 / 43.88 2.993 / 43.82 3.061 / 43.78 3.236 / 43.70
IV-I Impact of Noise Variance Estimation in Entropy Modeling
In our main FLLIC design, the clean-entropy estimation branch takes the noise standard deviation $\sigma_{N}$ as input to inform its prediction of $H(I)$. While this was initially studied under oracle conditions with ground-truth $\sigma_{N}$ provided, we further assess the impact of replacing this oracle information with either (i) the estimated $\widehat{\sigma}_{N}$ from the NLE or (ii) removing $\sigma_{N}$ altogether. Table IV reports the comparative results. Using the estimated $\widehat{\sigma}_{N}$ leads to only marginal bitrate overhead ($\leq 1.36\%$) compared to the ground-truth, confirming the effectiveness of the NLE. Omitting $\sigma_{N}$ entirely causes a larger bitrate increase (up to $7.18\%$), demonstrating that providing even an estimated noise level is beneficial. Nonetheless, FLLIC without $\sigma_{N}$ input still outperforms cascaded baselines, verifying its robustness.
TABLE IV: Impact of ground-truth vs. estimated vs. absent noise variance on rate–distortion performance. BPP / PSNR (dB).
$\sigma$ GT $\sigma_{N}$ Est. $\widehat{\sigma}_{N}$ No $\sigma_{N}$
1 3.009 / 48.91 3.050 / 48.90 ($+1.35\%$) 3.225 / 48.78 ($+7.18\%$)
2 2.959 / 45.71 2.990 / 45.70 ($+1.05\%$) 3.150 / 45.60 ($+6.45\%$)
3 2.802 / 43.88 2.840 / 43.85 ($+1.36\%$) 2.960 / 43.72 ($+5.64\%$)
IV-J Noise-Adaptive Extension of FLLIC
In the main experiments, FLLIC was trained with a separate model for each known noise level ($\sigma$) to assess its performance upper bound under controlled conditions. While this design facilitates fair benchmarking, it is not realistic for real-world scenarios where noise characteristics are often unknown or time-varying. To address this limitation, we propose a noise-adaptive variant of FLLIC that generalizes to unseen noise levels without requiring retraining.
The noise-adaptive FLLIC augments the original model with two lightweight components: (i) Noise-Level Estimator (NLE). A four-layer depthwise–separable CNN ($<\!0.15$M parameters) that predicts the global noise standard deviation $\widehat{\sigma}_{N}$ from the noisy input. The NLE is trained jointly with the codec, requiring no additional supervision; (ii) Conditional Modulation. The predicted $\widehat{\sigma}_{N}$ is injected into the entropy-estimation and quantization branches ($H(I)$, $Q_{y}$, $Q_{z}$) via FiLM-style affine modulation, enabling dynamic adjustment of coding strategies according to the estimated noise.
We train a single adaptive FLLIC model on mixed AWGN data ($\sigma=1\!-\!5$) and evaluate it on unseen noise levels ($\sigma=1.5,3.5,4$). The results in Table V show: (i) Compared to expert per-noise models, the adaptive model incurs only a $1.93\%$ average bitrate overhead under oracle $\sigma$; (ii) When relying on the NLE prediction, the additional penalty remains below $0.8\%$, verifying the estimator’s accuracy. Besides, the NLE adds only 3.6 ms latency per $512\!\times\!512$ image on RTX 4090, with no extra memory at decoding.
TABLE V: Evaluation of the noise-adaptive FLLIC variant. Reported: BPP / PSNR (dB) / Overhead.
$\sigma$ Expert FLLIC Adaptive (Oracle) Adaptive (Estimator)
$\sigma=1.5$ 2.980 / 47.20 3.035 / 47.22 ($+1.85\%$) 3.055 / 47.23 ($+2.52\%$)
$\sigma=3.5$ 2.810 / 43.10 2.865 / 43.12 ($+1.96\%$) 2.880 / 43.14 ($+2.49\%$)
$\sigma=4$ 2.780 / 42.30 2.835 / 42.33 ($+1.98\%$) 2.850 / 42.35 ($+2.52\%$)
Bitrate overhead $+1.93\%$ $<0.8\%$ vs Oracle